OpenStudy (anonymous):

What is y = x^2 -14x + 30 written in vertex form? y = (x - 7)^2 + 30 y = (x - 7)^2 - 19 y = (x - 14)^2 + 37 y = (x - 14)^2 + 49

OpenStudy (anonymous):

I know it's not y = (x - 7)^2 + 30. I'm leaning more towards y = (x - 7)^2 + 30, but I'd like to be sure :)

OpenStudy (anonymous):

y = x^2-14x +30 y = (x^2-14x+49)+30-49 y = (x-7)^2-19

OpenStudy (anonymous):

you should also take it look at the other problem you posted

OpenStudy (anonymous):

Thank you! (oh and it was supposed to say y=(x-7)^2-19 above but didn't copy paste it properly haha)

OpenStudy (anonymous):

No Problem